Absorption Statement

Absorption costing does not distinguish between variable and fixed costs. All manufacturing costs are included in the cost of goods sold.

Saxon, Inc.
Absorption Costing Income Statement
For the Year Ended December 31
Sales $1,125,000
Cost of goods sold:
  Cost of goods manufactured $840,000
  Ending inventory (210,000)
    Total cost of goods sold (630,000)
Gross profit $495,000
Selling and administrative expenses (275,000)
Operating income $220,000

Variable Statement

Under variable costing, the cost of goods manufactured includes only variable manufacturing costs. This type of income statement includes a computation of manufacturing margin.

Saxon, Inc.
Variable Costing Income Statement
For the Year Ended December 31
Sales $1,125,000
Variable cost of goods sold:
  Variable cost of goods manufactured $600,000
  Ending inventory (150,000)
    Total variable cost of goods sold (450,000)
Manufacturing margin $675,000
Variable selling and administrative expenses (210,000)
Contribution margin $465,000
Fixed costs:
  Fixed manufacturing costs $240,000
  Fixed selling and administrative expenses 65,000
    Total fixed costs (305,000)
Operating income $160,000

Method Comparison

Review the income statements on the Absorption Statement and Variable Statement, then complete the following table. The company’s sales price per unit is $75, and the number of units in ending inventory is 5,000. There was no beginning inventory.

Item Amount
Number of units sold
Variable sales and administrative cost per unit $
Number of units manufactured
Variable cost of goods manufactured per unit $
Fixed manufacturing cost per unit $
Answer #1

Calculate following

1) Number of unit sold = 1125000/75 = 15000 Unit

2) Variable sales and administrative cost per unit = 210000/15000 = 14 per unit

3) Number of unit manufactured = 15000+5000 = 20000 Units

4) Variable cost of goods manufactured per unit = 600000/20000 = 30 per unit

5) Fixed manufacturing cost per unit = (840000-600000)/20000 = 12 per unit